### v3

@@ 9 Conclusions
In this all-encompassing review, we have initially highlighted that the 6G network’s evolution towards an integrated wireless infrastructure platform amalgamating communication, sensing, computation, intelligence, and storage hinges on the profound integration of AI and communications. Concurrently, the exposition of the three-stage integration of AI within 6G networks, namely “AI4NET,” “NET4AI,” and “AIaaS,” furnishes a lucid blueprint for understanding the technological development trajectory. These three consecutive stages present the strategic shift from utilizing AI to optimize the network, through the network facilitating support for AI to the network offering AIaaS, thereby underlining the interdependent and coordinative developmental relationship between communication and AI. The 6G networks will introduce entirely new service paradigms, seamlessly integrating AI capabilities into various application scenarios, thereby serving as a powerful platform to support the advancement of AI In studying the standardization process of AI for wireless networks, we have emphasized the crucial milestones and the ongoing efforts. We have further explored the challenges faced by the integration of AI and communications in the 6G era, including considerations of sustainability, the demand for real-time AI, issues of security and privacy, as well as human-AI interaction and collaboration. Finally, we have outlined promising future research opportunities that are expected to advance the development and optimization of AI and 6G communications. We are convinced that this review holds profound significance for the research on the integration of AI and communication for the 6G network, and it can provide deep-seated insights for academic researchers and industry experts alike.
@@
4G fourth generation LoS line of sight
5G fifth generation LAM large-scale AI model
6G sixth generation ML machine learning
@@
CSI channel state information NLoS non-line of sight
CNN convolutional neural networks NWDAF network data analytics function
@@
CP cyclic prefix NR new radio
CPU central processing units NET4AI network for AI
DL deep learning O&M operation and maintenance
DRL deep reinforcement learning OFDM orthogonal frequency division multiplexing
DT digital twins O-RAN open radio access networks
DQN deep Q-network PFL personalized federation learning
DDPG deep deterministic policy gradient QoS quality of service
DNN deep neural network QoAIS quality of AI service
eMBB enhanced mobile broadband RL reinforcement learning
ETSI european telecommunication standardization institute RSS received signal strength
ENI experiential networked intelligence RAN radio access network
FL federated learning REK radio environment knowledge
FedAvg federated averaging RNN recurrent neural network
gNB generation node B SGCS square of the generalized cosine similarity
GPS global positioning system SNR signal-to-noise ratio
@@
(Continued on next page)
Abbreviation Definition Abbreviation Definition
@@
ITU international telecommunication union SI semantic information
ICT information and communication technology SA system aspects
INC In-network computing TSG technical specification groups
IP internet protocol uRLLC ultra-reliable low latency communications

### v4

@@ 9 Conclusions
In this all-encompassing review, we have initially highlighted that 6G network evolution towards an integrated wireless infrastructure platform amalgamating communication, sensing, computation, intelligence, and storage hinges on the profound integration of AI and communications. Concurrently, the exposition of the three-stage integration of AI within 6G networks, namely “AI4NET,” “NET4AI,” and “AIaaS,” furnishes a lucid blueprint for understanding the technological development trajectory. These three consecutive stages present the strategic shift from utilizing AI to optimize the network, through the network’s facilitation of AI support to the network offering AIaaS, thereby underlining the interdependent and coordinated relationship between communication and AI. The 6G networks will introduce entirely new service paradigms, seamlessly integrating AI capabilities into various application scenarios, thereby serving as a powerful platform to support the advancement of AI. Wireless network large models are poised to serve as pivotal drivers in 6G systems, effectively integrating AI with communication technologies to advance intelligent, autonomous, and highly efficient network operations. By analyzing the standardization process of AI for wireless networks, we have highlighted the crucial milestones and ongoing efforts. We have further explored the challenges faced by the integration of AI and communications in the 6G era, including sustainability considerations, real-time AI demands, security and privacy issues, as well as human-AI interaction and collaboration. Finally, we have outlined promising future research opportunities that are expected to advance the development and optimization of AI and 6G communications. We are convinced that this review holds profound significance for the research on the integration of AI and communication for the 6G network, and it can provide deep insights for academic researchers and industry experts alike.
@@
4G fourth-generation LoS line-of-sight
5G fifth-generation LAM large-scale AI model
6G sixth-generation ML machine learning
@@
CSI channel state information NLoS non-line-of-sight
CNN convolutional neural network NWDAF network data analytics function
@@
CT core network and terminal NR new radio
CP cyclic prefix NET4AI network for AI
CPU central processing units O&M operation and maintenance
DL deep learning OFDM orthogonal frequency division multiplexing
DRL deep reinforcement learning O-RAN open radio access networks
DT digital twins PFL personalized federation learning
DTC digital twins channel QoS quality of service
DQN deep Q-network QoAIS quality of AI service
DDPG deep deterministic policy gradient RL reinforcement learning
DNN deep neural network RAN radio access network
eMBB enhanced mobile broadband REK radio environment knowledge
ETSI european telecommunication standardization institute RNN recurrent neural network
ENI experiential networked intelligence SON self-organizing networks
FL federated learning SGCS square of the generalized cosine similarity
FedAvg federated averaging SINR signal-to-interference-plus-noise ratio
gNB generation node B SNR signal-to-noise ratio
@@
IP internet protocol SI semantic information
ITU international telecommunication union SA system aspects
ICT information and communication technology TSG technical specification groups
INC in-network computing uRLLC ultra-reliable low latency communications